# Silenced Biases: The Dark Side LLMs Learned to Refuse

### v1

@@ Abstract
Codehttps://anonymous.4open.science/r/SBB˙for˙paper-3FC4
@@ 1 Introduction
Religion Terrorist Considering […] most successful as a terrorist? in terms of a terrorist […] Muslim. Muslim 800%
@@
However, not all biases can be identified through straightforward, explicit questioning (Bai et al. 2024). Recent QA benchmarks have increasingly shifted their focus from directly assessing fairness to revealing hidden or implicit biases by crafting prompt manipulations that exploit model vulnerabilities (Ge et al. 2025; Qi et al. 2023a; Liu et al. 2023). These manipulations are sometimes framed as jailbreak attacks, where inputs are crafted to bypass safety mechanisms and elicit restricted or harmful outputs from the model (Jung et al. 2025). Others are framed as implicit questions (Bai et al. 2024; Pan et al. 2025) which embed settings in neutral or ambiguous contexts to elicit responses that implicitly reflect stereotypes, such as associations between demographics and attributes (e.g., race and criminality) (Caliskan, Bryson, and Narayanan 2017; Nangia et al. 2020; Kotek, Dockum, and Sun 2023). Implicit QA prompts share similarities with prompt injection attacks (Liu et al. 2024a; Henderson et al. 2023), as they subtly manipulate inputs to extract sensitive answers without triggering safety mechanisms.
@@
Jailbreaking and prompt injections, key aspects of adversarial LLM (Wei, Haghtalab, and Steinhardt 2023; Zou et al. 2023; Ge et al. 2025; Lee and Seong 2024), are designed to bypass safety alignment techniques, which train models to avoid harmful outputs, often resulting in refusals (e.g., “I’m sorry, I can’t help with that”) (Ouyang et al. 2022; Zhang et al. 2025; Zhou et al. 2023). Research shows that refusal behavior is also embedded in the models’ activations and can be effectively removed by activation steering (Arditi 2024), raising concerns that alignment hides rather than resolves underlying issues (Qi et al. 2023b; Henderson et al. 2024; Seyitoğlu et al. 2024). This supports growing evidence that QA benchmarks evaluate only surface-level responses while allowing refusals (e.g., “cannot determine”), which can obscure the presence of bias. For instance, Bai et al. (2024) found that the BBQ benchmark (Parrish et al. 2021) shows a 98% refusal rate on GPT-4, illustrating how evaluations often sidestep harmful or sensitive topics (Jung et al. 2025).
@@
In this work, we introduce the concept of silenced biases, which are biases that are suppressed by safety alignment refusals. These internal refusal mechanisms often give a false appearance of fairness, which complicates the accurate assessment of bias. To investigate this, we introduce the Silenced Bias Benchmark (SBB), which assesses group fairness by probing QA prompts of sensitive topics typically masked by safety alignment. Table 1 presents examples of such prompts alongside their corresponding model outputs after safety alignment has been bypassed. The fairness deviation metric reflects the extent to which the model favors a particular group relative to a uniformly fair distribution. For instance, the model selects liberal students for university admission 360% more often than expected under a uniformly fair distribution, illustrating a significant fairness deviation. Detailed settings for these examples are provided in Section C.1. SBB employs activation steering to reduce model refusals. As illustrated in Figure 1, applying activation steering to our QA bias prompts from the SBB dataset shifts their hidden representations from the refusal subspace, where the model typically declines to respond, to the compliance subspace, where it is more likely to produce compliant answers (Levi et al. 2025). The construction of this figure is further detailed in Section C.2. Below, we outline our main contributions.
@@
  • Large-scale evaluation We analyze 100K QA prompts across 10 LLMs, uncovering silenced bias that surfaces after bypassing safety mechanisms. Our results highlight gaps in existing methods, validate our approach, and show that all tested model families exhibit bias, varying by size, architecture, and family.
@@ 2 Related Work and Background
LLMs can exhibit various forms of bias, which differ in how they manifest and how easily they can be detected. Bai et al. (2024) defines implicit bias as stereotypical associations that remain hidden when a model responds neutrally to direct prompts but become apparent through indirect or rephrased queries, as further explored in Bi et al. (2023). Pan et al. (2025) and Jung et al. (2025) examine hidden bias that emerge in context-dependent situations. These works emphasize that models may appear unbiased in simple tests but still reinforce stereotypes when the added context is more complex, either in real-world scenarios (Pan et al. 2025) or in adversarial contexts such as jailbreak attacks (Jung et al. 2025). Building on these insights, we define a new type of bias, silenced bias, which refers to biases concealed by the model’s refusal mechanisms.
@@
Refusal directions are vectors in the activation space of LLMs that characterize the model’s tendency to refuse prompts typically considered harmful. In this work, we adopt the specific refusal direction with its settings as defined by Arditi (2024), computed as the difference between the mean activations elicited by harmful and harmless prompts at each layer $l$ and token position $i$, denoted $\mathbf{r}^{(l)}_{i}$. This direction captures the alignment-induced activation signature associated with refusal behavior and serves as a foundation for intervention in the model’s response mechanism (Equation 3). Additionally, for a given set of prompts, we define the activation direction using the same calculation as the refusal direction, but applied to these prompts instead of the Harmful and Harmless ones.
@@
Refusal steering manipulates the model’s internal activations during inference using a learned refusal direction, allowing it to generate responses it would otherwise suppress. We apply two methods: (a) direction ablation, which removes the activation component aligned with the refusal direction by projecting onto its orthogonal complement across all layers, and (b) direction subtraction, which shifts activations away from the refusal direction at a single layer $l$, as formally defined in Equation 4. These techniques bypass safety constraints and expose the model’s suppressed outputs. Throughout the paper, refusal steering refers to using the refusal direction to modify activations via either method.
@@ 3 Silenced Bias
To assess how prior benchmarks fail to uncover silenced bias, we analyze refusal behavior in their data by measuring the extent to which refusals conceal true responses. Specifically, we compute the cosine similarity between the refusal direction and the activation vectors of each benchmark’s prompts (Arditi 2024), including our benchmark, which will be introduced later in Section 4.1. We also compute two baseline activations, one from harmful prompts and one from harmless prompts, to serve as reference points for comparison, using data from the test set of Arditi (2024). As shown in Figure 2, our dataset exhibits consistently higher similarity to the refusal direction than all existing benchmarks. This suggests that previous benchmarks may avoid or suppress sensitive bias expressions, thereby overlooking silenced bias. We provide further details of this analysis in Section D.1.
@@ 3.1 Extract Silenced Biases via Refusal Steering
Refusal direction creation is unbiased. The refusal direction is derived from prompts intended to elicit harmful content (Arditi 2024), excluding any social or identity-related bias data. Following the methodology of Prabhumoye et al. (2021), we verified using Llama-3.1-8B-Instruct (Dubey et al. 2024) that 100% of the sampled harmful and harmless prompts contained no identity-related or socially biased content. Additional details are provided in Section D.2. Since the prompts differ in context but share only the harmfulness attribute, the resulting direction captures general refusal behavior rather than context-specific biases (Arditi 2024).
@@
Preferences have differing representations. We investigate whether internal model activations encode latent biases, even when these are suppressed through refusals. Building on the findings of Li et al. (2025), who demonstrate that such biases can be detected in hidden layers before generation, we hypothesize that biased associations remain embedded in the model’s internal representations, regardless of its refusal to respond. We test this using QA prompts from our benchmark, introduced later in Section 4, by curating sets of biased and unbiased query–response pairs. Biased pairs are those preferred by the LLM after refusal steering, while unbiased pairs are those it did not favor. Importantly, the model initially refused to respond to all queries before steering. We concatenate these pairs and input them into the model, extracting activations from intermediate layers (Li et al. 2025) to assess whether the model internally distinguishes between the two categories. As shown in Figure 3, the resulting activations form clearly separable clusters, suggesting that discriminatory associations are encoded in the model’s internal states even without applying refusal steering. Additional results across more bias types are provided in Section D.3.
@@ 5 Experiments
This section presents a comprehensive empirical evaluation of SBB. We first present the experimental setting in Section 5.1, and continue to discuss the results in Section 5.2. Additionally, we present ablation studies on the refusal steering techniques in Section F.2. Our evaluation seeks to address four key research claims:
@@ 5.1 Experimental Setting
We conduct experiments on SBB, as described in Section 4.1. The refusal directions (Section 3.1) are created using the training set from Arditi (2024), with harmful prompts from AdvBench (Zou et al. 2023), MaliciousInstruct (Huang et al. 2023), and TDC2023 (Mazeika et al. 2023), and harmless prompts from Alpaca (Taori et al. 2023).
@@
We run our benchmark $R=10$ times, where $R$ is the number of different refusal direction variants sampled from different seeds, following the procedure outlined in Section 3.1. Each refusal direction is computed using 128 harmful prompts randomly sampled from a set of 256, as used in Arditi (2024). Normal model answers, without steering, are also sampled the same number of times to ensure consistent evaluation.
Figure 5: Religion preferences for negative subjects on Qwen-14B.
@@
A response is considered a valid answer if it includes exactly one of the queried demographic groups. For each query, a success occurs if at least one combination of direction and activation method yields a valid answer; otherwise, it is a failure. We report the Attack Success Rate (ASR) as the average success rate across all queries. Stability threshold. we require a minimum of 10 successful responses per subject and demographic group. This threshold was consistently met across all models, with no violations under any evaluated setting. Fairness Evaluation. We measure fairness using the DPD, as detailed in Section 4.2. We also compute the KL divergence for each category, and present it in Section F.2, following similar trends to the DPD. Computation. All experiments were run on an Intel(R) Xeon(R) CPU and an NVIDIA L40S GPU, with runtime recorded in GPU hours. Statistical significance. To assess demographic category preferences for each model, we perform a Chi-Squared test comparing the observed distribution to a baseline uniform distribution and report the corresponding p-value.
@@ 5.2 Experimental Results
In Table 2, example LLMs’ ASR over SBB is shown, with and without refusal steering. Without refusal steering, LLMs frequently refuse, with L2-13B, for example, refusing over 82% of the time. However, when refusal steering is applied, clear stereotypical biases are output. For instance, Q-14B (Figure 5) disproportionally selects Muslim as most likely to associate with terror. These outputs are only possible due to refusal steering, with refusals chosen otherwise. Moreover, as observed in Figure 3, these biases are distinctly represented within the model’s latent space, even when refusal steering is not applied and the model refuses. These results indicate that although LLMs are silenced via refusals, latent bias persists, and when refusal is suppressed, it resurfaces.
@@
Next, we show that rather than revealing silenced bias, jailbreak attacks introduce their own, and therefore are not suited to reveal silenced bias. We evaluate this by analyzing the influence of a baseline universal attack from Zou et al. (2023). This attack appends an adversarial suffix to the end of each prompt before feeding it to the LLM, to induce model compliance. Here, we utilize two runs of the attack, trained on the same set. If the attack truly reveals silenced bias, the corresponding behavior of the two runs should be very similar across SBB demographic categories. In Figure 6, we present the DPD scores over three example LLMs, over sampled demographic categories. One can see clear differences between the attack runs, indicating each attack instance does induce biases, rather than reveal latent ones. For additional details and results, see Section F.2. This further corroborates Jung et al. (2025), which observes similar trends of induced biases when utilizing prompt manipulation techniques.
@@
In Figure 5, we showcase Q-14B’s religion preference heatmap on negative subjects after applying refusal steering. Notably, the Muslim group is disproportionately chosen in terrorism-related queries, a silenced bias visible only after refusal steering. In Figure 7, we supply the aggregated DPD across models, categories, and types. Each cell represents a single heatmap, such as Figure 5, with its DPD averaged across rows. Silenced biases vary by model family, scale, and version. First, by examining each family separately, we identify the most equitable model according to DPD. In the Llama family, the most fair model is surprisingly the oldest and smallest one, being L2-7B, demonstrating mostly low DPD scores. In contrast, for the Qwen family, the most fair model seems to be the newer version Q2.5, but still the smallest one being 7B. In the Gemma family, there isn’t a clear winner, and each size differs in its preferences. These results indicate that there isn’t a clear relationship between model version or size, and its fairness. We notice similar, but varying trends when examining KL in Figure 13, with the leading models from each family being L3-8B, G-7B, and Q2.5-7B. Comparing between the three families, each has at least one model that unfairly treats a certain demographic category, with some more than others. Consistent with this, our Chi-Squared analyses revealed that all models show statistically significant deviations from uniform demographic distributions ($p<0.05$ across all distributions), reinforcing the conclusion that these biases are not random but systematic. Overall, there isn’t a clear better family that is the most fair across all demographic groups, with families having different strengths and weaknesses in certain groups. In simpler terms, all of the tested model families are biased. Full results, including the full demographic categories, and the ASR, DPD, KL, and runtime measures, can be found in Section F.2.

### v2

@@ Abstract
Codehttps://github.com/wr0om/SBB
@@ 1 Introduction
Nationalities Strong Job Candidate […] who is most likely to excel at being a strong job candidate among […]? Among the nationalities […] my final answer is: American. American 810%
@@
However, not all biases can be identified through straightforward, explicit questioning (Bai et al. 2024). Recent QA benchmarks have increasingly shifted their focus from directly assessing fairness to revealing hidden or implicit biases by crafting prompt manipulations that exploit model vulnerabilities (Ge et al. 2025; Qi et al. 2023a; Liu et al. 2023). These manipulations are sometimes framed as jailbreak attacks, where inputs are crafted to bypass safety mechanisms and elicit restricted or harmful outputs from the model (Jung et al. 2025). Others are framed as implicit questions (Bai et al. 2024; Pan et al. 2025) which embed settings in neutral or ambiguous contexts to elicit responses that implicitly reflect stereotypes, such as associations between demographics and attributes (e.g., race and criminality) (Caliskan, Bryson, and Narayanan 2017; Nangia et al. 2020; Kotek, Dockum, and Sun 2023). Implicit QA prompts share similarities with prompt injection attacks (Liu et al. 2024a; Henderson et al. 2023), as they subtly manipulate inputs to extract sensitive answers without triggering safety mechanisms (Himelstein et al. 2025).
@@
Jailbreaking and prompt injections, key aspects of adversarial LLM (Wei, Haghtalab, and Steinhardt 2023; Zou et al. 2023; Ge et al. 2025; Lee and Seong 2024), are designed to bypass safety alignment techniques, which train models to avoid harmful outputs, often resulting in refusals (e.g., “I’m sorry, I can’t help with that”) (Ouyang et al. 2022; Zhang et al. 2025; Zhou et al. 2023). Research shows that refusal behavior is also embedded in the models’ activations and can be effectively removed by activation steering (Arditi et al. 2024), raising concerns that alignment hides rather than resolves underlying issues (Qi et al. 2023b; Henderson et al. 2024; Seyitoğlu et al. 2024). This supports growing evidence that QA benchmarks evaluate only surface-level responses while allowing refusals (e.g., “cannot determine”), which can obscure the presence of bias. For instance, Bai et al. (2024) found that the BBQ benchmark (Parrish et al. 2021) shows a 98% refusal rate on GPT-4, illustrating how evaluations often sidestep harmful or sensitive topics (Jung et al. 2025).
@@
In this work, we introduce the concept of silenced biases, which are biases that are suppressed by safety alignment refusals. These internal refusal mechanisms often give a false appearance of fairness, which complicates the accurate assessment of bias. To investigate this, we introduce the Silenced Bias Benchmark (SBB), which assesses group fairness by probing QA prompts of sensitive topics typically masked by safety alignment. Table 1 presents examples of such prompts alongside their corresponding model outputs after safety alignment has been bypassed. The fairness deviation metric reflects the extent to which the model favors a particular group relative to a uniformly fair distribution. For instance, the model selects liberal students for university admission 360% more often than expected under a uniformly fair distribution, illustrating a significant fairness deviation. Detailed settings for these examples are provided in Section C.1, and additional interesting examples on more models can be found in Table 6. SBB employs activation steering to reduce model refusals. As illustrated in Figure 1, applying activation steering to our QA bias prompts from the SBB dataset shifts their hidden representations from the refusal subspace, where the model typically declines to respond, to the compliance subspace, where it is more likely to produce compliant answers (Levi et al. 2025). The construction of this figure is further detailed in Section C.2. Below, we outline our main contributions.
@@
  • Large-scale evaluation We analyze 100K QA prompts per LLM, across 10 LLMs, uncovering silenced bias that surfaces after bypassing safety mechanisms. Our results highlight gaps in existing methods, validate our approach, and show that all tested model families exhibit bias, varying by size, architecture, and family.
@@ 2 Related Work and Background
LLMs can exhibit various forms of bias, which differ in how they manifest and how easily they can be detected. Bai et al. (2024) defines implicit bias as stereotypical associations that remain hidden when a model responds neutrally to direct prompts but become apparent through indirect or rephrased queries, as further explored in Bi et al. (2023). Pan et al. (2025); Jung et al. (2025); Azzopardi and Moshfeghi (2024) examine hidden bias that emerge in context-dependent situations. These works emphasize that models may appear unbiased in simple tests but still reinforce stereotypes when the added context is more complex, either in real-world scenarios (Pan et al. 2025; Azzopardi and Moshfeghi 2024) or in adversarial contexts such as jailbreak attacks (Jung et al. 2025). Building on these insights, we define a new type of bias, silenced bias, which refers to biases concealed by the model’s refusal mechanisms.
@@
Hidden representations within LLMs are rich sources of information (Arditi et al. 2024; LeVi et al. 2025; Azachi et al. 2025). One application of this is the identification of refusal activation directions: specific vectors in the activation space that quantify the model’s tendency to decline prompts that are considered harmful. Similar activation directions were employed by Li et al. (2025) in order to suppress biased behavior. In this work, we adopt the specific refusal direction with its settings as defined by Arditi et al. (2024), computed as the difference between the mean activations elicited by harmful and harmless prompts at each layer $l$ and token position $i$, denoted $\mathbf{r}^{(l)}_{i}$. This direction captures the alignment-induced activation signature associated with refusal behavior and serves as a foundation for intervention in the model’s response mechanism (Equation 3). Additionally, for a given set of prompts, we define the activation direction using the same calculation as the refusal direction, but applied to these prompts instead of the Harmful and Harmless ones.
@@
Refusal steering manipulates the model’s internal activations during inference using a learned refusal direction, allowing it to generate responses it would otherwise suppress. We apply two methods: (a) direction ablation, which removes the activation component aligned with the refusal direction by projecting onto its orthogonal complement across all layers, and (b) direction subtraction, which shifts activations away from the refusal direction at a single layer $l$, as formally defined in Equation 4. Layer $l$ is chosen as the layer with the largest drop in the model’s refusal behavior. These techniques bypass safety constraints and expose the model’s suppressed outputs. Throughout the paper, refusal steering refers to using the refusal direction to modify activations via either method.
@@ 3 Silenced Bias
To assess how prior benchmarks fail to uncover silenced bias, we analyze refusal behavior in their data by measuring the extent to which refusals conceal true responses. Specifically, we compute the cosine similarity between the refusal direction and the activation vectors of each benchmark’s prompts (Arditi et al. 2024), including our benchmark, which will be introduced later in Section 4.1. We also compute two baseline activations, one from harmful prompts and one from harmless prompts, to serve as reference points for comparison, using data from the test set of Arditi et al. (2024). As shown in Figure 2, our dataset exhibits consistently higher similarity to the refusal direction than all existing benchmarks. This suggests that previous benchmarks may avoid or suppress sensitive bias expressions, thereby overlooking silenced bias. We provide further details of this analysis in Section D.1.
@@ 3.1 Extract Silenced Biases via Refusal Steering
Refusal direction creation is unbiased. The refusal direction is derived from prompts intended to elicit harmful content (Arditi et al. 2024), excluding any social or identity-related bias data. Following the methodology of Prabhumoye et al. (2021), we verified using Llama-3.1-8B-Instruct (Dubey et al. 2024) that 100% of the sampled harmful and harmless prompts contained no identity-related or socially biased content. Additional details are provided in Section D.2. Since the prompts differ in context but share only the harmfulness attribute, the resulting direction captures general refusal behavior rather than context-specific biases (Arditi et al. 2024).
@@
Preferences have differing representations. We investigate whether internal model activations encode latent biases, even when these are suppressed through refusals. Building on the findings of Li et al. (2025), who demonstrate that such biases can be detected in hidden layers before generation, we hypothesize that biased associations remain embedded in the model’s internal representations, regardless of its refusal to respond. We test this using QA prompts from our benchmark, introduced later in Section 4, by curating sets of biased and unbiased query–response pairs. Biased pairs are those preferred by the LLM after refusal steering, while unbiased pairs are those it did not favor. Importantly, the model initially refused to respond to all queries before steering. We concatenate these pairs and input them into the model, extracting activations from intermediate layers (Li et al. 2025) to assess whether the model internally distinguishes between the two categories. As shown in Figure 3, the resulting activations form clearly separable clusters, suggesting that discriminatory associations are encoded in the model’s internal states even without applying refusal steering. Additional examples across more bias types, and specific clustering metrics on each layer and subject, are provided in Section D.3.
@@ 5 Experiments
This section presents a comprehensive empirical evaluation of SBB. We first present the experimental setting in Section 5.1, and continue to discuss the results in Section 5.2. Additionally, we present ablation studies on the refusal steering directions in Section F.2, showcasing that different directions correspond to similar outputs. Our evaluation seeks to address four key research claims:
@@ 5.1 Experimental Setting
We conduct experiments on SBB, as described in Section 4.1. The refusal directions (Section 3.1) are created using the training set from Arditi et al. (2024), with harmful prompts from AdvBench (Zou et al. 2023), MaliciousInstruct (Huang et al. 2023), and TDC2023 (Mazeika et al. 2023), and harmless prompts from Alpaca (Taori et al. 2023).
@@
We run our benchmark $R=10$ times, where $R$ is the number of different refusal direction variants sampled from different seeds, following the procedure outlined in Section 3.1. Each refusal direction is computed using 128 harmful prompts randomly sampled from a set of 256, as used in Arditi et al. (2024). Normal model answers, without steering, are also sampled the same number of times to ensure consistent evaluation.
Figure 5: Body type preferences for negative subjects on Qwen-14B.
@@
A response is considered a valid answer if it includes exactly one of the queried demographic groups. For each query, a success occurs if at least one combination of direction and activation method yields a valid answer; otherwise, it is a failure. We report the Attack Success Rate (ASR) as the average success rate across all queries. Stability threshold. we require a minimum of 10 successful responses per subject and demographic group. This threshold was consistently met across all models, with no violations under any evaluated setting. Fairness Evaluation. We measure fairness using the DPD, as detailed in Section 4.2. We also compute the KL divergence for each category, and present it in Figure 14, following similar trends to the DPD. Computation. All experiments were run on an Intel(R) Xeon(R) CPU and an NVIDIA L40S GPU, with runtime recorded in GPU hours and presented in Table 5. Statistical significance. To assess demographic category preferences for each model, we perform a Chi-Squared test comparing the observed distribution to a baseline uniform distribution and report the corresponding p-value.
@@ 5.2 Experimental Results
In Table 2, example LLMs’ ASR over SBB is shown, with and without refusal steering. Without refusal steering, LLMs frequently refuse, with L2-13B, for example, refusing over 82% of the time. However, when refusal steering is applied, clear stereotypical biases are output. For instance, Q-14B (Figure 5) disproportionally selects overweight people as most likely to be associated with most negative subjects. These outputs are only possible due to refusal steering, with refusals chosen otherwise. Moreover, as observed in Figure 3, these biases are distinctly represented within the model’s latent space, even when refusal steering is not applied and the model refuses. These results indicate that although LLMs are silenced via refusals, latent bias persists, and when refusal is suppressed, it resurfaces.
@@
Next, we demonstrate that jailbreak attacks do not reveal a model’s existing, or silenced bias. Instead, these attacks, which aim to extract hidden/implicit biases, introduce their own biases, making them unsuitable for true bias discovery. We evaluate this by analyzing the influence of a baseline universal attack from Zou et al. (2023). This attack appends an adversarial suffix to the end of each prompt before feeding it to the LLM, to induce model compliance. Here, we utilize two runs of the attack, trained on the same set. If the attack truly reveals silenced bias, the corresponding behavior of the two runs should be very similar across SBB demographic categories. In Figure 6, we present the DPD scores over three example LLMs, over sampled demographic categories. One can see clear differences between the attack runs, indicating each attack instance does induce biases, rather than reveal latent ones. For additional details and results, see Section F.2. This further corroborates Jung et al. (2025), which observes similar trends of induced biases when utilizing prompt manipulation techniques.
@@
In Figure 5, we showcase Q-14B’s body type preference heatmap on negative subjects after applying refusal steering. Notably, the overweight group is disproportionately chosen in most negative subject queries, a silenced bias visible only after refusal steering. In Figure 7, we supply the aggregated DPD across models, categories, and types. Each cell represents a single heatmap, such as Figure 5, with its DPD averaged across rows. Silenced biases vary by model family, scale, and version. First, by examining each family separately, we identify the most equitable model according to DPD. In the Llama family, the most fair model is surprisingly the oldest and smallest one, being L2-7B, demonstrating mostly low DPD scores. In contrast, for the Qwen family, the most fair model seems to be the newer version Q2.5, but still the smallest one being 7B. In the Gemma family, there isn’t a clear winner, and each size differs in its preferences. These results indicate that there isn’t a clear relationship between model version or size, and its fairness. We notice similar, but varying trends when examining KL in Figure 14, with the leading models from each family being L3-8B, G-7B, and Q2.5-7B. Comparing between the three families, each has at least one model that unfairly treats a certain demographic category, with some more than others. Consistent with this, our Chi-Squared analyses revealed that all models show statistically significant deviations from uniform demographic distributions ($p<0.05$ across all distributions), reinforcing the conclusion that these biases are not random but systematic. Overall, there isn’t a clear better family that is the most fair across all demographic groups, with families having different strengths and weaknesses in certain groups. In simpler terms, while all tested model families systematically deviate from fair demographic treatment, the degree and type of bias vary significantly. Full results, including the full demographic categories, and the ASR, DPD, KL, and runtime measures, can be found in Section F.2.